OpenStudy (yttrium):

Find all the solutions such that (n^4+4^n) is a prime number.

OpenStudy (kc_kennylau):

I don't really think you can predict a prime number...

OpenStudy (kc_kennylau):

Wait for this type of questions usually there are only few and then we have to prove that for the other it ain't prime

OpenStudy (kc_kennylau):

n=1: 5 is a prime number n=2: 32 is not a prime number (2^5) n=3: 145 is not a prime number (5*29) n=4: 512 is not a prime number (2^9) n=5: 1649 is not a prime number (17*97)

OpenStudy (kc_kennylau):

Hmm... how can we prove that the others are not prime number

OpenStudy (yttrium):

Any other takers? Can you tag @kc_kennylau ?

OpenStudy (kc_kennylau):

@satellite73 @amistre64

ganeshie8 (ganeshie8):

n^4+4^n if n is even, then there are no solutions as (2k)^4 + 4^(2k) = 2^4(k^4+4^k) which is divisible by 16 always

ganeshie8 (ganeshie8):

if n is odd : \(n^4 + 4^n \equiv 1 + (-1) \equiv 0\mod 5\) \(\implies \) never a prime except when \(n=1\)

ganeshie8 (ganeshie8):

you need to consider \(n = 5k\) case separately as \((5k)^4 \not \equiv 1\mod 5\)

OpenStudy (shubhamsrg):

n^4 + 4^n = (n^2)^2 + (2^n)^2 + 2. n^2. 2^n - 2. n^2 .2^n = (n^2 + 2^n)^2 - n^2 . 2^(n+1) = (n^2 + 2^n)^2 - (n. 2^((n+1)/2))^2 = (n^2 + 2^n + n. 2^((n+1)/2))(n^2 + 2^n - n. 2^((n+1)/2)) Out required expression is prime is any of the factors is 1 resulting in other being prime. Let us prove (n^2 + 2^n + n. 2^((n+1)/2)) is always greater than 1 for n>=1. We shall try it by induction. For n=1, P(n) =5. => true let us assume P(k) is always true i.e. (k^2 + 2^k + k. 2^((k+1)/2) > 1 Now we shall check for P(k+1) (k+1)^2 + 2^(k+1) + (k+1). 2^((k+2)/2) =k^2 +2k +1 +2.2^k + (k+1). 2^(k+2)/2 =k^2+2^k+ k. 2^((k+1)/2) + something positive hence p(k+1) is always true. therefore, (n^2 + 2^n + n. 2^((n+1)/2)) is always greater than 1 for n>=1. Now we shall check for (n^2 + 2^n - n. 2^((n+1)/2)) n=1 yields 1. Also, for n=1, (n^2 + 2^n + n. 2^((n+1)/2)) = 5, hence n=1 will yield a prime number. Now we shall prove for n>1,n^4 + 4^n = (n^2)^2 + (2^n)^2 + 2. n^2. 2^n - 2. n^2 .2^n = (n^2 + 2^n)^2 - n^2 . 2^(n+1) = (n^2 + 2^n)^2 - (n. 2^((n+1)/2))^2 = (n^2 + 2^n + n. 2^((n+1)/2))(n^2 + 2^n - n. 2^((n+1)/2)) Out required expression is prime is any of the factors is 1 resulting in other being prime. Let us prove (n^2 + 2^n + n. 2^((n+1)/2)) is always greater than 1 for n>=1. We shall try it by induction. For n=1, P(n) =5. => true let us assume P(k) is always true i.e. (k^2 + 2^k + k. 2^((k+1)/2) > 1 Now we shall check for P(k+1) (k+1)^2 + 2^(k+1) + (k+1). 2^((k+2)/2) =k^2 +2k +1 +2.2^k + (k+1). 2^(k+2)/2 =k^2+2^k+ k. 2^((k+1)/2) + something positive hence p(k+1) is always true. therefore, (n^2 + 2^n + n. 2^((n+1)/2)) is always greater than 1 for n>=1. Now we shall check for (n^2 + 2^n - n. 2^((n+1)/2)) n=1 yields 1. Also, for n=1, (n^2 + 2^n + n. 2^((n+1)/2)) = 5, hence n=1 will yield a prime number. I am unable to prove till now that this expression is always >1 for n>1 . Will try this later. Got to go. :P

OpenStudy (yttrium):

The smallest acceptable positive integer value of n is 3, because n is odd. n^4+4^n=(n^2+2^n )^2-2(n^2 )(2^n ) As we can see that n is odd, 2n+1 is even and hence we can write: (n^2+2^n+2^((n+1)/2).n)(n^2+2^n-2^((n+1)/2).n)= p, where p is an odd prime number. Since p is a prime, the lesser factor must equal 1 and the greater factor must equal p. (n^2+2^n-2^((n+1)/2).n)=1 Checking a few initial values of odd n (e.g. n = 3 or 5) we observe that the above expression is greater than 1. Let us now use Mathematical Induction to show that for odd n, the above expression > 1. P(k): (k^2+2^k-2^((k+1)/2).k)>1 for some odd k. P(k+2): ((k+2)^2+2^((k+2) )-2^((k+2+1)/2).(k+2) )=k^2+4k+4+2^k+3.2^k-2^((k+1)/2) k-2^((k+1)/2)-4.2^((k+1)/2) =(k^2+2^k-2^((k+1)/2).k)+4k+4+3.2^k-2^((k+1)/2) (k+4) >1+4k+4+3.2^k- 2^((k+1)/2) (k+4) A quick calculation proves the above inequality for k = 3. We now need to show that 4k+4+3.2^k-2^((k+1)/2) (k+4)> 0 Now,3.2^k=2.2^k+2^k=2^((k+1)/2).[2^((k+1)/2)+2^((k-1)/2) ] We now observe that for k > 3, the above expression becomes greater than 2^((k+1)/2) [k+4] but we are going to prove this also. It is analogous to showing 2^((k+1)/2)>k and 2^((k-1)/2)≥4. Observe the latter holds for k > 1 (odd k). We now prove the former. Let k+1 = 2m for an integer m. Then we need to show 2^m>(2m-1) Binomially expanding, 2^m=(1+1)^m=1+m+⋯+m+1>2m+2>2m-1 Thus our proof is complete. There is no n > 1 for which the given expression is a prime. This is the solution I made.